40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 17 18:44:29 2013, Masayuki, Update, LSC, LSC calibration screen calibration_screen.png
    Reply  Fri Sep 20 22:49:10 2013, Masayuki, Update, LSC, LSC calibration screen plot.pdf
       Reply  Wed Sep 25 08:11:01 2013, Masayuki, Summary, LSC, LSC calibration screen plot.pdfScreenshot.pngScreenshot-1.pngfringe.pdf
          Reply  Fri Sep 27 13:55:11 2013, Masayuki, Update, LSC, LSC calibration screen 
             Reply  Fri Sep 27 21:01:50 2013, Masayuki, Update, LSC, LSC calibration screen 
                Reply  Mon Sep 30 11:33:15 2013, rana, Update, LSC, LSC calibration screen 
Message ID: 9158     Entry time: Wed Sep 25 08:11:01 2013     In reply to: 9149     Reply to this: 9169
Author: Masayuki 
Type: Summary 
Category: LSC 
Subject: LSC calibration screen 

 

  The real time calibration system is not correct in high frequency.

The attachment are the plot of two free running noise. Blue curve is the plot of noise calibrated with OLTF. Green one is the just fft analysed signal of the real time calibration system output. You can see the ripple in high frequency region in green curve. That is because the anti-aliasing filter and digital anti-aliasing filter. I assume the sensitivity of MI as constant but Rana mentioned that we should take these filters into account.

modeled OLTF and sensitivity H
 I put the AA filter and DAA filter effects into matlab calibration script. The attachment 2 is the modeled sensitivity of the MICH. You can find each filter properties in  elog#8555 (analog AA filter) and in elog#3961. I estimate the H gain by measuring the fringe. The attachment 3 is the plot of fringe and I averaged with green points. The actual number is 3.48e7 count/m

attachment 2: the sensitivitiy of MICH

Screenshot.png

 attachment 3: fringe of the MICH

fringe.pdf

I modeled OLTF with this H and the fitted into the measurement data. That is in attachment 4. In this OLTF I also included the DAI filter and AI filter, and ' sample and hold circuit' of DAC TF . These are  mentioned in two references. Additionally I added the time delay 309.6 us.  Yuta mentioned that in C1SUS has 125us time delay. In MICH control we have also C1LSC , so I think this time delay is reasonable. I compensated the error signal with these OLTF and MICH sensitivity.

attachment 4: OLTF of the MICH control

Screenshot-1.png

You can see that the ripple is gone in blue curve and after 5 kHz the curve is flat.

Next step

I'm trying to put the inverted AA filter and DAA filter in C1CAL_INCV servo. But the ploblem is the difference of sampling frequency, so I couldn't fix yet. One possibility is putting approximated filter. I hope I will find some good way to design these filters.

 

 

Other thing

I esitimated the FPMI noise propagated from the residual noise of IR PDH control of both ARMS. I will summarize and write these staff in this afternoon.

Attachment 1: plot.pdf  68 kB  | Hide | Hide all
plot.pdf
ELOG V3.1.3-